OpenStudy (anonymous):

If y varies directly as x, and y is 20 when x is 4, what is the constant of variation for this relation?

OpenStudy (danjs):

solve y = k*X for k ,, when x = 4 and y=20

OpenStudy (danjs):

k is the constant of variation

OpenStudy (anonymous):

80?

OpenStudy (danjs):

20 = k * 4 k = 20/4 = ?

OpenStudy (anonymous):

5

OpenStudy (danjs):

yes